The surface area of a rectangular box with a
square base is 54 square inches, and is given
by
A = 2x2 + 4xh
Find x when h = 3 inches.

Answer by ankor@dixie-net.com(22740)

The surface area of a rectangular box with a square base is 54 square inches,
and is given by:
A = 2x^2 + 4xh
2x^2 + 4xh = 54
:
Find x when h = 3 inches.
2x^2 + 4x(3) = 54
2x^2 + 12x - 54 = 0
Simplify divide by 2
x^2 + 6x - 27 = 0
Factors to
(x+9)(x-3) = 0
Positive solution
x = 3
:
:
Check this, find the surface area with x=3,h=3
A = 2(3^2) + 4*3*(3)
A = 2(9) + 36
A = 18 + 36
A = 54